OpenStudy (anonymous):

lim n-infity 3n+1/2n+1=3/2

OpenStudy (john_es):

This is a limit that can be evaluated in various ways, \[\lim_{n\rightarrow\infty}\frac{3n+1}{2n}=\lim_{n\rightarrow\infty}\frac{3n/n+1/n}{2n/n}=3/2\] We have assumed that in the limit proposed, 1/n reaches 0.